# MACE-OFF: Short Range Transferable Machine Learning Force Fields for Organic Molecules

### v4

@@ Abstract
preprint: AIP/123-QED**footnotetext: These authors contributed equally.
@@ II.1 The MACE architecture
×$\qquad\qquad\times Y^{m_{1}}_{l_{1}}(\bm{\hat{r}}_{ij})h^{(t)}_{j,kl_{2}m_{2}}$ (2)
@@ III.4.1 Water structure and dynamics
Figure 7: Water density as a function of temperature for MACE-OFF23(M), and the extended cutoff MACE-OFF24(M) model.
@@ III.4.2 Ala3 Free Energy Surface
Figure 8: Free energy surfaces of Ala3 in explicit solvent. (left) AMBER14SB/TIP3P and (right) MACE-OFF24(M).
@@ Acknowledgements
Contributions
DPK conceived the project, curated the training data, trained and evaluated the models, ran parts of the torsion drive tests, the molecular crystal tests, fitted the dipole and polarisability models for water and paracetamol bulk systems, and wrote the initial draft. JHM and DPK implemented the OpenMM interface. JHM trained and evaluated the models, curated the datasets, performed experiments on molecular liquids, peptides and proteins, benchmarked the model performance in OpenMM and wrote the initial draft. JTH implemented the MACE-torsion drive interfaces, and ran the DFT and force field torsion drives. VK fitted the MACE-OFF23-$\mu$, performed the MD and PIMD simulations of water and molecular crystals and produced their IR and Raman spectra. YP performed the error analysis for the MACE-OFF23-$\mu$ model and performed MD and PIMD simulations and calculated the IR spectrum of paracetamol molecule. IB and NJB implemented the accelerated custom kernels, and the cuda_mace library is maintained by NJB. WCW implemented the MACE-LAMMPS interface, wrote the pure C++ and Kokkos implementations, and benchmarked the model performance in LAMMPS. IBM developed the intermolecular error analysis and evaluated the models. DJC and GC supervised the project. All authors discussed the results and edited the final manuscript.

### v5

@@ Abstract
preprint: AIP/123-QED22footnotetext: These authors contributed equally.11footnotetext: Corresponding author: jhm72@cam.ac.uk
@@ II.1 The MACE architecture
$\qquad\qquad\times Y^{m_{1}}_{l_{1}}(\bm{\hat{r}}_{ij})h^{(t)}_{j,kl_{2}m_{2}}$ (2)
@@ III.4.1 Water structure and dynamics
Figure 7: Water density as a function of temperature. Comparison between MACE-OFF23(M), and the extended cutoff MACE-OFF24(M) model.
@@ III.4.2 Ala3 Free Energy Surface
Figure 8: Free energy surfaces of Ala3 in explicit solvent. AMBER14SB/TIP3P (top) and MACE-OFF24(M) (bottom).
@@ Acknowledgements
Elements of this work are adapted from a PhD thesis.
@@
Supplementary Information
The supplementary information contains the following sections:
  • Training details for the MACE-OFF models, including filtering of the SPICE dataset
  • Detailed breakdown of test set errors, including intramolecular force errors
  • Details of the torsion scans
  • Details of lattice enthalpy calculations
  • Crystal structure geometries
  • Details of condensed phase simulations, including error tables
  • Details of biological simulations
  • Simulation details and results for dipole models
  • Computational performance in LAMMPS